OpenStudy (anonymous):

Which functions have a positive rate of change and which have a negative rate of change?

OpenStudy (anonymous):

Column A 1. 2x – 3y = 9 2. y = −4x + 5 3. x y −1 −5 1 1 3 7 4. x y 1 −2 2 −6 3 −10

OpenStudy (anonymous):

Column B A. positive B. negative

OpenStudy (anonymous):

Do you have any way of graphing these?

OpenStudy (anonymous):

nope sorry

OpenStudy (igreen):

Rate of change is also known as slope.

OpenStudy (anonymous):

Okay then.. one sec.

OpenStudy (igreen):

2x – 3y = 9 To find the slope of this, change it into slope-intercept form(y = mx + b): 2x - 3y = 9 Subtract 2x to both sides: -3y = -2x + 9 Divide -3 to both sides: y = 2/3x + 3 Now it's in the form of y = mx + b, where m = slope. So here the slope or average rate of change is POSITIVE 2/3.

OpenStudy (igreen):

y = -4x + 5 Do you know what the slope is here? @RainbowBubbles01

OpenStudy (anonymous):

hold on

OpenStudy (igreen):

It's already in the form of y = mx + b.

OpenStudy (anonymous):

no i dont -_- i have been so busy this week i am so burnt out ,luckily today is friday

OpenStudy (igreen):

y = mx + b m = slope y = -4x + b What's the slope? Hint: the number next to \(x\).

OpenStudy (anonymous):

-4

OpenStudy (igreen):

Yep, and -4 is negative so y = -4x + 5 has a negative slope.

OpenStudy (anonymous):

sooo....

OpenStudy (igreen):

Now for #3, we can take two points from the table and plug them into the slope formula. Let's take (1, 1) and (3, 7): \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{7-1}{3-1}\) Subtract: \(m = \dfrac{6}{2}\) Can you divide that? @RainbowBubbles01

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait divide 6/2?

OpenStudy (igreen):

Yes.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

.

OpenStudy (igreen):

Yep, and that's positive. So so far we have: 1 - Positive 2 - Negative 3 - Positive Now we do the same thing with #4: Let's take (2, -6) and (3, -10): \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-10+6}{3-2}\) Add & Subtract: \(m = \dfrac{-4}{1}\) Can you divide that? @RainbowBubbles01

OpenStudy (anonymous):

-4

OpenStudy (igreen):

Yep, and that's negative so: Question 1 - Positive Question 2 - Negative Question 3 - Positive Question 4 - Negative

OpenStudy (anonymous):

thank you thank you thank you, can you help with 4 more?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Sure, just close this one first and open a new one.